A brass rod is 2m long at a certain temperature . what is it's lenght for a temperature rise of 100k if the expression vity of brass is 1.8*10^_6 k^7


1
Expert's answer
2022-01-17T09:59:10-0500

Given:

"l_0=\rm 2\: m"

"\Delta T=\rm 100\: K"

"\alpha=\rm 18.8*10^{-6}\: K^{-1}"


The equation of linear temperature expansion of a length is given by

"l=l_0(1+\alpha\Delta T)"

So, the new length

"l=\rm 2\: m*(1+18.8*10^{-6}*100)=2.0038\: m"